OpenStudy (anonymous):

Solve the problem. Find (f + g)(-4) when f(x) = x + 4 and g(x) = x - 7.

jimthompson5910 (jim_thompson5910):

(f+g)(x) = f(x)+g(x) (f+g)(x) = (x + 4) + (x - 7) (f+g)(x) = x + 4 + x - 7 (f+g)(x) = 2x - 3 (f+g)(-4) = 2(-4) - 3 (f+g)(-4) = -8 - 3 (f+g)(-4) = -11

OpenStudy (anonymous):

argh you beat me, but yeah that's right